OpenStudy (anonymous):

what is 19 aproximated to the nearest tenth

OpenStudy (anonymous):

is 19 closer to 10 or 20?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

there you go.

OpenStudy (anonymous):

\[whatis \sqrt{19} aproximated \to the nearest tenth\]

OpenStudy (anonymous):

the two perfect squares next to 19 are \[\sqrt{16}\] and \[\sqrt{25}\]